OpenStudy (anonymous):

Find the equation of a line which passes through point (-7, 1) and is parallel to the line formed by the equation y = 6x + 8. Express it in slope-intercept form. y = - 1 x + 43 y = 6x + 43 y = 6x - 43 y=(1/6)x+43

OpenStudy (chaise):

y = 6x+8 If the line is parralel, then they have the same gradient, with a different y intercept (8). So, if the second line runs through (-7,1) we can substitute these values into the "y = mx+b" with m = 6 (same gradient as before) y = mx+b 1 = 6x-7+b 1 = -42 + b 43 = b Your y intercept of this equation is 43. We can now put this all together to form the equation: y = mx+b y = 6x+43. Does this make sense? Do you need further explanation?